Richie’s utility function is given by U=5X+2Y.

Draw Richie’s indifference curve when U=10.

What is the marginal rate of substitution of X for Y?

What is the marginal rate of substitution when X=1 and Y=5? When X=2 and Y=2.5?
Answer #1

Richies utiliti utility function # when U= 5x+27 U = 10 10 = 5x +2y Y У 3 2 IC, X 2 3 4 x =-5 = 22.5 2 * MRS my = -MUs -Əv (